8: Convergence of summation (n!(n+1)!)/(3n)!

9: Radius of convergence of summation n=0 to infinity (n^3)x^3n/n^4+1


1
Expert's answer
2021-03-24T13:49:41-0400

8: Let us study the convergence of summation "\sum_{n=1}^{\infty}\frac{n!(n+1)!}{(3n)!}". Let us use the D'Alembert's rule.


Taking into account that


"\lim\limits_{n\to \infty}(\frac{(n+1)!(n+2)!}{(3(n+1))!}:\frac{n!(n+1)!}{(3n)!})=
\lim\limits_{n\to \infty}(\frac{(n+1)n!(n+2)(n+1)!}{(3n+3)!}\cdot\frac{(3n)!}{n!(n+1)!})=
\lim\limits_{n\to \infty}\frac{(n+1)(n+2)}{(3n+3)(3n+2)(3n+1)}=
\lim\limits_{n\to \infty}\frac{(1+\frac{1}{n})(1+\frac{2}{n})}{(3+\frac{3}{n})(3+\frac{2}{n})(3n+1)}=0<1"


we conclude that "\sum_{n=1}^{\infty}\frac{n!(n+1)!}{(3n)!}" is convergent.


9: Let us find the radius of convergence of summation "\sum_{ n=0}^{\infty}\frac{ n^3x^{3n}}{n^4+1}". Let us use the D'Alembert's rule.


Since "\lim\limits_{n\to \infty}|\frac{ (n+1)^3x^{3(n+1)}}{(n+1)^4+1}:\frac{ n^3x^{3n}}{n^4+1}|=
|x|^3\cdot\lim\limits_{n\to \infty}\frac{ (n+1)^3(n^4+1)}{((n+1)^4+1)n^3}=
|x|^3\cdot\lim\limits_{n\to \infty}\frac{ (1+\frac{1}{n})^3(1+\frac{1}{n^4})}{(1+\frac{1}{n})^4+\frac{1}{n^4}}=|x|^3\cdot 1=|x|^3<1"imply "|x|<1", we conclude that the radius "R" of convergence of summation "\sum_{ n=0}^{\infty}\frac{ n^3x^{3n}}{n^4+1}" is "R=1."